# Sparse–Smooth Spatially Varying Coefficient Quantile Regression

### v1

@@ 3 Estimation and Algorithms
$\mathcal{G}_{h}(\bm{\theta}_{P},\bm{\Theta}_{L})=\sum_{i=1}^{n}M_{h}\Big(r_{i}(\bm{\theta}_{P},\bm{\Theta}_{L})\Big)+\lambda_{2}\sum_{j=1}^{p}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},\qquad\mathcal{R}(\bm{\Theta}_{L})=\lambda_{1}\sum_{j=1}^{p}w_{j}\|\bm{\delta}_{j}\|_{2}.$
@@
$\nabla_{\bm{\alpha}}\mathcal{G}_{h}=-\bm{Z}^{\top}\bm{g},\quad\nabla_{\bm{\beta}_{G}}\mathcal{G}_{h}=-\bm{X}^{\top}\bm{g},\quad\nabla_{\bm{\delta}_{j}}\mathcal{G}_{h}=-\bm{X}_{\odot j}^{\top}\bm{g}+2\lambda_{2}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},$
@@
$\bm{\alpha}\leftarrow\bm{\alpha}-t\nabla_{\bm{\alpha}}\mathcal{G}_{h},\qquad\bm{\beta}_{G}\leftarrow\bm{\beta}_{G}-t\nabla_{\bm{\beta}_{G}}\mathcal{G}_{h},$
@@
$\varepsilon_{\mathrm{pri}}=\sqrt{n}\varepsilon_{\mathrm{abs}}+\varepsilon_{\mathrm{rel}}\max\{\|\bm{s}\|_{2},\ \|\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}\|_{2}\},$
$\varepsilon_{\mathrm{dual}}=\sqrt{n}\varepsilon_{\mathrm{abs}}+\varepsilon_{\mathrm{rel}}\max\{\|\rho_{s}(\bm{s}-\bm{s}^{\mathrm{prev}})\|_{2},\ (\sum_{j}\|\rho_{z}(\bm{z}_{j}-\bm{z}_{j}^{\mathrm{prev}})\|_{2}^{2})^{1/2}\}.$
@@ 4 Asymptotic Theory
$\mathcal{L}_{n}(\bm{\theta}_{P},\bm{\Theta}_{L})=\sum_{i=1}^{n}\rho_{\tau}\Big(Y_{i}-\bm{Z}_{i}^{\top}\bm{\alpha}-\bm{X}_{i}^{\top}\bm{\beta}_{G}-\sum_{j=1}^{p}X_{ij}\delta_{j}(u_{i})\Big)+\sum_{j=1}^{p}\lambda_{1n}w_{j,n}\|\bm{\delta}_{j}\|_{2}+\sum_{j=1}^{p}\lambda_{2n}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},$
@@
Let $(\bm{\alpha}_{0},\bm{\beta}_{0})$ denote the true parameters in the sense of (3). Write $\beta_{0,j}(u)=\beta_{G0,j}+\delta_{0,j}(u)$ with degree-weight centered $\delta_{0,j}(\cdot)$ and define
$\bm{\delta}_{j0}=\big(\delta_{0,j}(u_{1}),\dots,\delta_{0,j}(u_{n})\big)^{\top},\qquad\mathcal{A}_{0}=\{j:\bm{\delta}_{j0}=\bm{0}\},\ \ \mathcal{A}_{0}^{c}=\{1,\dots,p\}\setminus\mathcal{A}_{0}.$
@@
$\bm{M}=\mathrm{E}\big[f_{\varepsilon\mid Z,X,U}(0)\bm{G}_{i}\bm{G}_{i}^{\top}\big],\qquad\bm{V}=\tau(1-\tau)\mathrm{E}\big[\bm{G}_{i}\bm{G}_{i}^{\top}\big].$
@@
    $\Pr\big(\hat{\bm{\delta}}_{j}=\bm{0}\ \text{for all }j\in\mathcal{A}_{0}\big)\to 1,\qquad\Pr\big(\hat{\bm{\delta}}_{j}\neq\bm{0}\ \text{for all }j\in\mathcal{A}_{0}^{c}\big)\to 1.$
@@
Proof sketch. A localized quadratic expansion of the empirical check loss in the direction of $\bm{\delta}_{j}$, combined with the Laplacian quadratic, yields a linear approximation whose stochastic fluctuation in the degree-weight centered subspace is of order $(n\lambda_{2n})^{-1/2}$. The deterministic bias is of order $\lambda_{2n}\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}$. Selection follows by a primal–dual witness argument together with the bound $\|\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\bm{r}_{0})\|_{2}=O_{p}(\sqrt{n})$ under bounded covariates.
@@
Sketch. By Cauchy–Schwarz and boundedness,
$\lesssim\sqrt{\frac{1}{n}\sum_{i=1}^{n}\Big(\sum_{j}X_{ij}e_{j,i}\Big)^{2}}\ \lesssim\ \sqrt{\frac{1}{n}\sum_{i,j}e_{j,i}^{2}}\ =\ O_{p}\left(\sqrt{\frac{1}{n}\sum_{j}\mathrm{MSE}_{\delta,j}}\right)=\ O_{p}(n^{-1/4})\to 0,$
using Theorem 3(i). This yields a negligible remainder in the Bahadur expansion of $\hat{\bm{\theta}}_{P}$.
@@ 5 Simulation Study
$Z_{i0}\equiv 1,\quad Z_{i1}\sim\mathcal{N}(0,1),\quad Z_{i2}\sim\mathcal{N}(0,1),\quad X_{ij}\sim\mathcal{N}(0,1)\ (j=1,\dots,4),$
@@
$\bm{\alpha}_{0}=(3,-1.0,1.5)^{\top},\qquad\bm{\beta}_{G0}=(5,0,2.5,0)^{\top},$
@@ 5.2 Monte Carlo Assessment
In order to objectively evaluate the performance of the proposed SS–SVCQR model, we compare it with five representative competitors. (i) Global quantile regression (QR) uses a purely global specification and thus ignores spatial nonstationarity. (ii) Standard geographically weighted regression (GWR) treats all covariates as spatially varying. (iii) Mixed GWR assumes that the local and global covariates are known a priori and fit a partially varying model; this can be regarded as an oracle benchmark for methods that must learn the local/global structure. (iv) Quantile generalized additive models (QGAM) allow for flexible nonlinear effects but do not directly target spatially varying coefficients. (v) The classical spatially varying coefficient (SVC) model is fitted by least squares without quantile weighting. These methods span the spectrum from purely global to fully spatially varying specifications and from mean to quantile regression, and they are chosen to assess three aspects: (a) accuracy of global and varying coefficient estimation, (b) ability to recover the true local versus global covariates, and (c) predictive performance under a range of error distributions. We report the detailed results of the Monte Carlo simulation in the six tables below. [^1] In Tables 16, the values in parentheses represent the standard deviation corresponding to that metric.
@@
Table 2: Standard_GWR
@@
Table 3: Mixed_GWR
@@
Table 4: QGAM
@@
Table 5: Classic_SVC
@@
Table 6: $S$-SVCQR
@@
Tables 16 summarise Monte Carlo means and standard deviations of the above metrics over $R=500$ replicates in Scenario 2. For the purely global quantile regression (QR) in Table 1, the global parameter error $\mathrm{PE}_{\theta}$ and deviation MSEs are substantially larger than those of the spatial competitors whenever spatial nonstationarity is present. Moreover, QR classifies all covariates as global, leading to sensitivity essentially equal to zero and specificity equal to one, and its predictive performance as measured by $\mathrm{CL}_{\mathrm{test}}$ is clearly inferior once the error distribution departs from the Gaussian case.
Standard GWR and mixed GWR (Tables 2 and 3) reduce $\mathrm{PE}_{\theta}$ and the deviation MSEs relative to QR, but they do so by treating almost all effects as spatially varying. As a consequence, sensitivity is essentially one while specificity is close to zero across all error distributions, indicating severe overfitting of local structure. QGAM (Table 4) shows a similar pattern: it achieves smaller $\mathrm{PE}_{\theta}$ and deviation MSEs than QR, but again classifies nearly all covariates as local, with sensitivity near one and specificity near zero. The classical SVC model in Table 5 behaves analogously, and under heavy-tailed Cauchy errors its parameter error and deviation MSEs can become extremely large despite reasonably small check loss in some scenarios.
In contrast, the proposed SS–SVCQR estimator in Table 6 delivers uniformly small global parameter errors and deviation MSEs across all error settings. With the larger penalty used here, SS–SVCQR achieves sensitivity essentially equal to one and specificity close to one for normal, heteroscedastic, $t_{3}$, and contaminated-normal errors, and only slightly below one for ALD and Cauchy errors. Thus SS–SVCQR is the only method that simultaneously recovers both truly local covariates $\{X_{1},X_{3}\}$ and truly global covariates $\{X_{2},X_{4}\}$ with high accuracy. Its predictive check loss $\mathrm{CL}_{\mathrm{test}}$ is uniformly lower than, or comparable to, that of the best competitor for light- and moderately heavy-tailed errors, and remains stable even under Cauchy errors, where several fully spatial methods suffer from dramatic loss of efficiency. These results confirm that SS–SVCQR provides near-oracle estimation of global coefficients, accurate recovery of local structure, and competitive predictive performance over a wide range of error distributions in Scenario 2.

### v2

@@ 3 Estimation and Algorithms
$\mathcal{G}_{h}(\bm{\theta}_{P},\bm{\Theta}_{L})=\sum_{i=1}^{n}M_{h}\Big(r_{i}(\bm{\theta}_{P},\bm{\Theta}_{L})\Big)+\lambda_{2}\sum_{j=1}^{p}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},\quad\mathcal{R}(\bm{\Theta}_{L})=\lambda_{1}\sum_{j=1}^{p}w_{j}\|\bm{\delta}_{j}\|_{2}.$
@@
$\nabla_{\bm{\alpha}}\mathcal{G}_{h}$ $=-\bm{Z}^{\top}\bm{g},$
$\nabla_{\bm{\beta}_{G}}\mathcal{G}_{h}$ $=-\bm{X}^{\top}\bm{g},$
$\nabla_{\bm{\delta}_{j}}\mathcal{G}_{h}$ $=-\bm{X}_{\odot j}^{\top}\bm{g}+2\lambda_{2}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},$
@@
$\bm{\alpha}\leftarrow\bm{\alpha}-t\nabla_{\bm{\alpha}}\mathcal{G}_{h},\quad\bm{\beta}_{G}\leftarrow\bm{\beta}_{G}-t\nabla_{\bm{\beta}_{G}}\mathcal{G}_{h},$
@@
$\varepsilon_{\mathrm{pri}}=\sqrt{n}\varepsilon_{\mathrm{abs}}+\varepsilon_{\mathrm{rel}}\max\left\{\|\bm{s}\|_{2},\left\|\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}\right\|_{2}\right\},$
$\varepsilon_{\mathrm{dual}}=\sqrt{n}\varepsilon_{\mathrm{abs}}+\varepsilon_{\mathrm{rel}}\max\left\{\|\rho_{s}(\bm{s}-\bm{s}^{\mathrm{prev}})\|_{2},\left(\sum_{j}\left\|\rho_{z}\left(\bm{z}_{j}-\bm{z}_{j}^{\mathrm{prev}}\right)\right\|_{2}^{2}\right)^{1/2}\right\}.$
@@ 4 Asymptotic Theory
$\mathcal{L}_{n}(\bm{\theta}_{P},\bm{\Theta}_{L})=\sum_{i=1}^{n}\rho_{\tau}\Big(Y_{i}-\bm{Z}_{i}^{\top}\bm{\alpha}-\bm{X}_{i}^{\top}\bm{\beta}_{G}-\sum_{j=1}^{p}X_{ij}\delta_{j}(u_{i})\Big)+\sum_{j=1}^{p}\lambda_{1n}w_{j,n}\|\bm{\delta}_{j}\|_{2}$
$+\sum_{j=1}^{p}\lambda_{2n}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},$
@@
Let $(\bm{\alpha}_{0},\bm{\beta}_{0})$ denote the true parameters in the sense of (3)[FanLi2001]. Write $\beta_{0,j}(u)=\beta_{G0,j}+\delta_{0,j}(u)$ with degree-weight centered $\delta_{0,j}(\cdot)$ and define
$\bm{\delta}_{j0}=\big(\delta_{0,j}(u_{1}),\dots,\delta_{0,j}(u_{n})\big)^{\top},\,\mathcal{A}_{0}=\{j:\bm{\delta}_{j0}=\bm{0}\},\,\mathcal{A}_{0}^{c}=\{1,\dots,p\}\setminus\mathcal{A}_{0}.$
@@
Intuitively, Assumption 1 requires that locations be well spread over a compact region and that the kk–NN graph does not degenerate as $n$ grows.
@@
$\bm{M}=\mathrm{E}\big[f_{\varepsilon\mid Z,X,U}(0)\bm{G}_{i}\bm{G}_{i}^{\top}\big],\quad\bm{V}=\tau(1-\tau)\mathrm{E}\big[\bm{G}_{i}\bm{G}_{i}^{\top}\big].$
@@
    $\Pr\big(\hat{\bm{\delta}}_{j}=\bm{0}\ \text{for all }j\in\mathcal{A}_{0}\big)\rightarrow 1,\quad\Pr\big(\hat{\bm{\delta}}_{j}\neq\bm{0}\ \text{for all }j\in\mathcal{A}_{0}^{c}\big)\rightarrow 1.$
@@ 5 Simulation Study
$Z_{i0}\equiv 1,\quad Z_{i1}\sim\mathcal{N}(0,1),\quad Z_{i2}\sim\mathcal{N}(0,1),\quad X_{ij}\sim\mathcal{N}(0,1),j=1,\dots,4,$
@@
$\bm{\alpha}_{0}=(3,-1.0,1.5)^{\top},\quad\bm{\beta}_{G0}=(5,0,2.5,0)^{\top},$
@@ 5.2 Monte Carlo Assessment
In order to objectively evaluate the performance of the proposed SS–SVCQR model, we compare it with five representative competitors. (i) Global quantile regression (QR) uses a purely global specification and thus ignores spatial nonstationarity[quantregPackage]. (ii) Standard geographically weighted regression (GWR) treats all covariates as spatially varying[GWmodelPackage]. (iii) Mixed GWR assumes that the local and global covariates are known a priori and fit a partially varying model; this can be regarded as an oracle benchmark for methods that must learn the local/global structure. (iv) Quantile generalized additive models (QGAM) allow flexible nonlinear effects but do not directly target spatially varying coefficients[Fasiolo2020]. (v) The classical spatially varying coefficient (SVC) model is fitted by least squares without quantile weighting[Hastie1993, Gelfand2003SVC]. These methods span the spectrum from purely global to fully spatially varying specifications and from mean to quantile regression, and they are chosen to assess three aspects: (a) accuracy of global and varying coefficient estimation, (b) ability to recover the true local versus global covariates, and (c) predictive performance under a range of error distributions. We report the detailed results of the Monte Carlo simulation in the six tables below. [^1] In Tables 16, the values in parentheses represent the standard deviation corresponding to that metric.
@@
Table 2: Monte Carlo result for geographically weighted regression
@@
Table 3: Monte Carlo result for mixed geographically weighted regression
@@
Table 4: Monte Carlo result for quantile generalized additive models
@@
Table 5: Monte Carlo result for spatially varying coefficient model
@@
Table 6: Monte Carlo result for $S$-SVCQR
@@
Tables 16 summarize Monte Carlo means and standard deviations of the above metrics in $R=100$ replicates in Scenario 2. For the purely global quantile regression (QR) in Table 1, the global parameter error $\mathrm{PE}_{\theta}$ and the deviation MSEs are substantially larger than those of the spatial competitors whenever there is spatial nonstationarity. Moreover, QR classifies all covariates as global, leading to a sensitivity essentially equal to zero and a specificity equal to one, and its predictive performance, measured by $\mathrm{CL}_{\mathrm{test}}$, is clearly inferior once the error distribution departs from the Gaussian case.
Standard GWR and mixed GWR (Tables 2 and 3) reduce $\mathrm{PE}_{\theta}$ and the deviation MSEs relative to QR, but do so by treating almost all effects as spatially varying. As a consequence, sensitivity is essentially one, while specificity is close to zero across all error distributions, indicating severe overfitting of local structure. QGAM (Table 4) shows a similar pattern: achieves smaller $\mathrm{PE}_{\theta}$ and deviation MSEs than QR, but again classifies nearly all covariates as local, with sensitivity near one and specificity near zero. The classical SVC model in Table 5 behaves analogously, and under heavy-tailed Cauchy errors, its parameter error and deviation MSEs can become extremely large despite reasonably small check loss in some scenarios.
In contrast, the proposed SS–SVCQR estimator in Table 6 delivers uniformly small global parameter errors and deviation MSEs across all error settings. With the larger penalty used here, SS–SVCQR achieves sensitivity essentially equal to one and specificity close to one for normal, heteroscedastic, $t_{3}$, and contaminated-normal errors, and only slightly below one for ALD and Cauchy errors. Thus, SS–SVCQR is the only method that simultaneously recovers both truly local covariates $\{X_{1},X_{3}\}$ and truly global covariates $\{X_{2},X_{4}\}$ with high precision. Its predictive check loss $\mathrm{CL}_{\mathrm{test}}$ is uniformly lower than, or comparable to, that of the best competitor for light- and moderately heavy-tailed errors, and remains stable even under Cauchy errors, where several fully spatial methods suffer from dramatic loss of efficiency. These results suggest that SS–SVCQR achieves near-oracle estimation of global coefficients, reliably recovers local structure, and maintains competitive predictive performance over a wide range of error distributions in Scenario 2.